OpenStudy (horsegal244):

An oil company fills 1 over 10 of a tank in 1 over 5 hour. At this rate, which expression can be used to determine how long will it take for the tank to fill completely?

OpenStudy (horsegal244):

@misty1212 @mathstudent55 @brunomcv

OpenStudy (anonymous):

1 over 10 (1/10) will be 0.1 which is 10% 1 over 5 (1/5) will be 0.5 which is 50% 10% in a 50% hour, it fills the tank 10% in 30 minutes. To get 100% multiply by 10, so in 300 Minutes it fills the tank. 300 minutes is 5 hours.

OpenStudy (anonymous):

You got it?

OpenStudy (horsegal244):

so 300 is the answer?

OpenStudy (anonymous):

In minutes, yes, in hours it's 5 To look more simple just answer 5 hours.

OpenStudy (horsegal244):

these are the answers

OpenStudy (horsegal244):

@brunomcv?

OpenStudy (anonymous):

Hold on, let me review this

OpenStudy (anonymous):

There's the error, 1 over 5 hours is 20% of an hour, my bad.

OpenStudy (horsegal244):

k

OpenStudy (anonymous):

1 over 10 (1/10) will be 0.1 which is 10% 1 over 5 (1/5) will be 0.2 which is 20% 10% in a 20% hour, it fills the tank 10% in 12 minutes. To get 100% multiply by 10, so in 120 Minutes it fills the tank. 120 minutes or 2 hours. According to ur available answers, it's the first one, (1/5)(10)

OpenStudy (horsegal244):

ok i have 1 more i need help with..

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

The rate is 1/10 of a tank in 1/5 hour. \(\dfrac{\frac{1}{10} ~tank}{\frac{1}{5} hour} \) We want a full tank, so we need to multiply the numerator by 10 to have 1 tank. \(=\dfrac{\frac{1}{10} ~tank}{\frac{1}{5} hour} \times \dfrac{10}{10} \) \(=\dfrac{\frac{1}{10} \times 10~tank}{\frac{1}{5} \times 10~hour}\) \(= \dfrac{1~tank}{2~hours} \)